$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155.
-----------------
The answer has to be one of the following but which?
-----------------
$1,000
$3,000
$4,000

Found 2 solutions by stanbon, Vladdroid:
Answer by stanbon(75887)

$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155.
-----------------
Equations:
x + y = 5000
3x + 3.5y = 15500
------
30x + 30y = 150000
30x + 35y = 155000
-----------------------
5y = 5000
----
y = $1000.00 (amt invested at 3.5%
---
x = $4000.00 (amt. invested at 3%
=====================================
Cheers,
Stan H.
=================

The answer has to be one of the following but which?
-----------------
$1,000
$3,000
$4,000

Answer by Vladdroid(91)

The answer is $4000.00.
Thank you